# Exploring the hardness of the ionizing radiation with the infrared softness diagram

### v1

@@ 4.1 A softness parameter in the mid-IR valid for AGN
Given the much harder SED assumed to be the cause of the ionization of the elements in the NLR of AGN, as compared to massive stars (see Fig. 1), we first explored to what extent the above-defined can be useful in tracing the hardness of the radiation field in these objects.
@@
To do so, we inspected the sequences of photoionization models described in the above subsection and obtained the values for the log parameter as a function of and log , represented in Fig. 2. In the left panels of the figure, we show the variation of log predicted by the models as a function of (upper panel) and log (lower panel) for two different values of metallicity (i.e., 12+log(O/H) = 9.0, 8.1). The models shown in the Figure assume a stopping criterion corresponding to a fraction of free electrons of 98%. Considering that higher values of correspond to harder SEDs, as can easily be verified in Fig. 1, no clear correlation can be seen between the value of , defined for its use in star-forming galaxies, and the hardness of the radiation field in AGN. This is due to the relatively low ionization potential of the lines involved in , which mostly probe the continuum of the big blue bump below 3 Ryd and therefore respond poorly to the power-law continuum at higher energies.
@@
Nonetheless, we can take advantage of the fact that in the mid-IR range we can measure higher ionization levels of the same species, which can help us extend the energetic range of the incident SED that we can assess using the available emission lines (e.g. Spinoglio et al., 2015; Fernández-Ontiveros et al., 2016). This is more helpful in distinguishing possible differences in the power-law. As an alternative measurable in the mid- to far-IR spectral range, we define:
@@
This modification allows us to largely extend the energetic range evaluated by the parameter, covering wavelengths where the possible differences in the power-law index defining the incident radiation field in AGN can be properly distinguished. Note that this parameter can also be defined using alternative equivalent transitions for the same ions in a similar spectral range, as in the case for [Ne v] at 24.3 m, and [O iii] at 88 m, which can replace the transitions defined in the above expression.
@@
In the middle panels of Fig. 2, we show the variations of this new parameter as a function of and log . Notice that we only represent the parameter when the models predict a certain emission of all involved lines, so it cannot be defined for certain low and for which the models do not predict any emission of the high-excitation lines. As can be seen, the parameter now clearly traces the hardness of the power-law, as it has lower values for higher values of , consistent with the original optical softness parameter defined by Vílchez & Pagel (1988), which decreases for harder radiation fields. Another important characteristic of this new parameter is its very low dependence on metallicity, given that the assumed AGN SEDs have not got any variation as a function of it, added to the relatively much lower dependence of the mid-IR emission-lines emissivities on electron temperature, as compared to optical lines.
@@
On the contrary, one important caveat of this new form of the softness parameter is the fact that the far-IR [O iii], either at 52 m or 88 m, are not covered at low redshift using the available facilities, once the SOFIA observatory has been retired, limiting the number of objects for which an analysis based on this parameter can be carried out. An alternative could be to replace this line using another emission line at a shorter wavelengths, more easily covered by mid-IR observations, corresponding to a similar energy range. One possibility could be the use of the sulphur emission lines of [S iv] at 10.5 m and [S iii] at 18.7 m, covering a very similar energy range (Fig. 1), although with a possible additional uncertainty coming from variations in the S/O abundance ratio (e.g. Pérez-Díaz et al., 2024). In this case, the alternative form of the softness parameter would be:
@@
As in the previous case, similar expressions can be explored using the [S iii] line at 33.5 m instead. In the right panels of Fig. 2 we show the dependence of this parameter on and log , where a very similar behaviour to that obtained for the version of the parameter based on [O iii] is seen, confirming its potential use to trace the hardness of the ionizing SED in AGN.
@@
In Fig. 3, we compare the two above-defined parameters represented in two dimensions under the form of diagrams, in order to compare in the same space the observations described in Section 2, represented as density points, with the predictions from some of the photoionization models described in Section 3. As a first evident impression, the use of the mid-IR lines of [S iii] and [S iv] replacing the far-IR [O iii] results very convenient, given the scarcity of points observed in that regime. In contrast, the IDEOS catalog allows us to explore a significantly larger sample. As support of this, in our overall sample of 484 objects, we only found 97 presenting a measurement of the far-IR [O iii] line, either at 52 m or 88 m, while [S iii]+[S iv], either with [S iii] at 18.7 m or 33.5 m, can be measured in 480 objects of the compiled sample.
@@
As it can be seen, although the sequences of larger cover the region of the diagram corresponding to high values for both ratios, they consistently show lower values (i.e. the lines of slope 1 in the diagram), while the models with lower are situated in agreement with the expected higher values of . Regarding , the values go in the same direction as already observed in the version of the softness diagram studied for star-forming regions: higher values of are situated in the lower left part of each sequence (i.e. corresponding to low values of Ne23Ne5 and O3O4), while lower values of are situated in the upper right part (i.e. for high values of Ne23Ne5 and O3O4). Both diagrams hence show that they could in principle be used to determine both and .
@@
Regarding the equivalent model sequences calculated without the assumption of any dust grains, in agreement with several authors who do not consider them in order to match the observations of certain UV lines (e.g. Nagao et al. 2006; Dors et al. 2019). However, in this case, although with a slight difference for lower values of , this does not seem to lead to results with differences larger than owing to other factors such as geometry, as can also been confirmed in the subsections below, probably motivated by the much lower dependence of the IR lines on dust absorption. This also underlines again the advantage of using IR lines for this kind of analysis, reducing the uncertainty due to extinction correction in the used emission lines.
4.2 The Hii-Chi-mistry-Teff-IR code for AGN
The use of the softness parameter based on IR lines, as defined in equations 2 and 3, or using alternative ratios of emission-line fluxes involving low- and high-excitation states of ionization, can serve as a proxy for the hardness of the incident radiation. However, as discussed in Pérez-Montero et al. (2019b) for the optical version of this parameter based on [O ii], [O iii], [S ii], and [S iii], the parameter presents additional dependencies on other functional parameters that can lead to incorrect interpretations for objects presenting different values but also with very different metallicities or excitation levels. A similar behavior was found and studied in Paper I when the corresponding version of the parameter for the mid-IR was examined.
@@ 4.2 The Hii-Chi-mistry-Teff-IR code for AGN
Similarly, although the mean of the different model sequences correlate with their , there are also deviations as a function of , as it can be seen in Fig.3, and in lesser term on , as it can be seen in Fig.2, that must be analyzed in order to extract a more precise estimation of the hardness of the radiation field. The code is thus a tool that helps to reduce these additional dependencies to be properly used before any derivation of the hardness of the incident radiation field. Therefore, we describe here the adaptation of the code HCm-Teff-IR to include the emission lines predicted by the photoionization model grids shown in Section 3. Additionally, we discuss the appropriate emission-line ratios required to construct the parameter under its several forms, adapted to the supplied observational information.
The code works in such a way that, once the parameters of the models that will be used (i.e. , dust) are chosen, it first interpolates the tables for the corresponding metallicity, which is treated as input by the code, expressed as the total oxygen abundance. As in previous versions, if the metallicity is not given as input, the code considers all possible values, implying an enhanced uncertainty.
The code then begins an iterative process along all the grid of models to create a -weighted distribution of and log , whose mean and standard deviation are the found solutions. Notice that this is not the classical Bayesian approach in which a unique model with the least resulting is chosen, as this last procedure would be very inefficient to figure out the real uncertainty of the result. The corresponding values are obtained by the code as the sum of the quadratic differences between the observed emission-line ratios and those predicted by the models, which can be calculated with the emission-line fluxes given as input. For this version of the code, we ruled out in the grids those models corresponding to and log , because these do not follow any clear trend useful to derive these parameters from the chosen emission-line ratios, as can be seen in Fig. 2.
@@
The code admits as input the reddening-corrected fluxes in arbitrary units of the emission lines of the ratios involved in the several forms of the parameter, including444Additionally, other emission-lines such as [Ar ii] 7.0 m + [Ar iii] 8.9 m and/or 21.8 m are also accepted by the code when they are given in relation to other high-excitation lines as [Ar v] 7.9 m or 13.1 m. We did not find any object in the compiled samples with these lines, but their interest does not present any doubt considering their potential detection in many targets observed by JWST in the near future. :
@@
Specifically for the case of the mid-IR lines in this study, we did not apply any extinction correction due to the relatively low extinction values involved. Additionally, the narrow-line emission in AGN is produced by an extended and diffuse gas region, which is les affected by the nuclear dust obscuration. The code only considers an emission-line flux for the calculation when it can be used in combination with the emission-line flux for a complementary ionization stage (e.g., [Ne v] is only used if, in addition, [Ne ii] and [Ne iii] are also given). In case more than one line is given for a certain ion (e.g., [Ne v], [S iii]), the code uses all forms of the corresponding ratios with equal weight. Therefore, it is up to the user to decide which line(s) to include if not all are to be used.
@@ 5.1 The distribution of for AGN
First, we calculated the parameters as defined for AGN in equations 2 and 3 for the compiled sample. The corresponding distributions for these values are represented in the histograms shown in Fig. 4. Note that for these calculations we chose the [O iii] line at 88 m instead of that at 52 m, as this allows us to calculate the parameter in a larger number of objects in our sample. A similar consideration was taken for the version of the parameter shown in the lower panel, for which we used the [S iii] line at 18.7 m instead of that at 33.5 m. Similarly, we used the [Ne v] line at 14.3 m.
@@
To verify the hypothesis that the studied points can indeed be represented by a double-peaked distribution, we applied a Gaussian Mixture Model (GMM), assigning each point to a different component. Afterwards, to confirm the bimodality found by the GMM, we applied a Shapiro-Wilk normality test independently to each of the two resulting distributions. In each case, the null hypothesis is that the objects do not correspond to a single Gaussian distribution, which is rejected if the resulting p-value is below a 5% threshold. Note that the uncertainties in the line ratios were not taken into account for the normality test. Nevertheless, this would have reduce the weight of uncertain measurements, which typically show larger deviations.
For both representations of the parameter (i.e. using either [O iii] or [S iii]+[S iv]), we verified that all p-values were below the rejection limit, confirming that each one of the two independent distributions found by the GMM are normal and hence that the whole resulting distribution is indeed bimodal. While the overlap region between the two Gaussian components found by the GMM introduces some uncertainty in individual classifications, the statistical trends remain robust as the fraction of galaxies with ambiguous membership is small and does not significantly impact the results.
We also show in Table LABEL:disp_eta the values obtained from this analysis with the mean and standard deviation resulting for each one of the two distributions. The mean values are also shown in Fig. 3. This result could thus indicate that two families of AGN with a different hardness could co-exist. Nevertheless, given the dependencies shown by both parameters for AGN (i.e. on and ), a more precise analysis is required to confirm this scenario.
@@ 5.2 Results from HCm-Teff-IR
We applied the HCm-Teff-IR code to our compiled sample of AGN with measured emission lines in the mid and far-IR spectral range, in order to break the possible degeneracy of the parameter. In particular, we aim to explore whether the observed bimodality is still recovered when we translate the information from the emission lines into the derived parameters and log , being the former a parametrization of the hardness of the incident radiation field in AGN, and the latter representing the excitation state of the gas.
@@
The use of the code also allows for a consistent comparison of results for objects presenting various emission-line combinations across different spectral ranges, thereby increasing the number of objects in the analysis. In this regard, the code admits as input any of the transition lines from [Ne v], [S iii], or [O iii], as described in Section 4.2. The only condition to provide a solution is that the Ne23Ne5 emission-line ratio, along with either O3O4 or S34O4, is available555Additionally, the code admits the emission-line ratio Ar23Ar5, although this is not available for any of the objects compiled for this work..
@@
In Table LABEL:disp we present the results for the objects for which the code finds a solution for , explored in the range from -2.0 to -1.0 and log from -4.0 to -1.0. The table shows the mean values and the standard deviations of the corresponding distributions under different assumptions adopted for the models, including the change in the selected stopping criterion. This value is varied from 2% of free electrons, which in practice implies a totally radiation-bounded geometry, to values 98% and 99.9%, which are those for which most of the sample is covered in the corresponding softness diagrams, as seen in Fig. 3. Results for models computed with or without dust grains mixed with the gas, for each value of , are also given. We did not find any noticeable difference neither in the results nor in their uncertainties for those objects for which a previous determination of the metal content was used by the code, underlining that metallicity is not a crucial factor to derive the hardness or the excitation using IR lines for AGN.
@@
From an inspection of these average results, it can be seen that the inclusion of dust in the models does not result in a noticeable difference in the outcomes. This is not unexpected, given the very small difference between the model predictions shown by the sequences in Fig. 3. On the contrary, the assumption of different stopping values implies that tends to be higher, on average, for decreasing . This is consistent with the fact that models with = 2% tend to have higher values. However, as discussed in previous sections, this implies that a non-negligible fraction of the studied sample lies outside the model grid, so it is preferable to use the values derived from the code when matter-bounded geometry is assumed, as suggested by previous works (e.g. Meléndez et al., 2011; Pérez-Montero et al., 2019a).
@@
Regarding the differences obtained for the two studied sub-samples (i.e. one for the objects with [O iii] and another for [S iii]+[S iv], which it is essentially the IDEOS sample), there are noticeable differences, particularly for the values, which are significantly higher in the case of IDEOS, with averages lower than -2.0. Nevertheless, Fig. 5 shows that the overall ranges in and log values derived are similar in both cases. Thus, the discrepancy observed could be ascribed to the different sample selection criteria in the two cases. In this regard, objects with [O iii] detections from Herschel/PACS were likely selected among the brightest nearest Seyfert nuclei to guarantee the detection of the faintest far-IR lines. A selection bias toward more luminous objects is likely associated with higher accretion rates.
@@
In addition to the average values, it is also important to analyze the shapes of the parameter distributions to probe how the bimodality found in is interpreted by the code. Fig. 5 shows the results obtained for the models with dust grains assuming = 99.9%. The left panel shows the subsample using the O3O4 ratio, either with [O iii] 52m or 88m, while the right panel corresponds to S34O4, using either [S iii] at 18m or 33m. The different symbols in the vs. log diagram correspond to the optical or IR spectral classes indicated in Véron-Cetty & Véron (2010) and Spoon et al. (2022), respectively. See Section 5.3 for a discussion on the distribution of different AGN types on this diagram. As in the previous analysis for , the number of objects in this second diagram is much larger (480 against 97), as it involves most of the galaxies in the IDEOS catalog. Note that the total number of objects with a successful determination of and log is 484 (Table LABEL:class), which differs from the number of values in Table LABEL:disp_eta. While is defined only for the 88 m [O iii] line (or alternatively the 18.7 m [S iii] line), the code provides a solution for all the objects with detected [Ne ii], [Ne iii], [Ne v] (14.3 or 24.3 m), [O iv] and either [O iii] (52 or 88 m), or [S iv] and [S iii] (18.7 or 33.5 m). The most important aspect, however, is that the bimodal distribution of the points remains for the two variables derived by the code, that is, for both and log . Besides, the bimodal nature of the two variables appears to be correlated, such that objects with a higher exhibit lower log and vice versa.
@@
Moreover, models help us to realize that the bimodality simultaneously appears for both the radiation field hardness and the excitation. Indeed, the correlation coefficient between these two variables is, on average across the different model assumptions, -0.95. In addition, this correlation can be easily confirmed by quantifying the relative number of objects simultaneously showing a higher (Hard Radiation, HR) and a lower log (Low Excitation, LE), or vice versa, i.e. those with a Soft Radiation field (SR) and a High Excitation (HE), which are the combinations to most of the galaxies apparently belong. As derived from the values shown in Table LABEL:class, this calculation yields that a 91.5% are SRHE or HRLE, confirming that objects with a harder radiation field also have a lower proportion of ionizing photons compared to the number of particles in the gas phase. Taking the weighted mean values obtained in Table LABEL:disp, we can thus define two main AGN families: one with average -1.7 and log -1.5 (SRHE), and another with -1.3 and log -2.4 (HRLE).

### v2

@@ 4.1 A softness parameter in the mid-IR valid for AGNs
Given the much harder SED assumed to be the cause of the ionization of the elements in the NLR of AGNs, as compared to massive stars (see Fig. 1), we first explored to what extent the above-defined can be useful in tracing the hardness of the radiation field in these objects.
@@
To do so, we inspected the sequences of photoionization models described in the above subsection and obtained the values for the log parameter as a function of and log , represented in Fig. 2. In the left panels of the figure, we show the variation of log predicted by the models as a function of (upper panel) and log (lower panel) for two different values of metallicity (i.e., 12+log(O/H) = 9.0, 8.1). The models shown in the figure assume a stopping criterion corresponding to a fraction of free electrons of 98%. Considering that higher values of correspond to harder SEDs, as can easily be verified in Fig. 1, no clear correlation can be seen between the value of , defined for its use in star-forming galaxies, and the hardness of the radiation field in AGNs. This is due to the relatively low ionization potential of the lines involved in , which mostly probe the continuum of the big blue bump below 3 Ryd and therefore respond poorly to the power-law continuum at higher energies.
@@
Nonetheless, we can take advantage of the fact that in the mid-IR range we can measure higher ionization levels of the same species, which can help us extend the energetic range of the incident SED that we can assess using the available emission lines (e.g., Spinoglio et al., 2015; Fernández-Ontiveros et al., 2016). This is more helpful in distinguishing possible differences in the power-law. As an alternative measurable in the mid- to far-IR spectral range, we defined
@@
This modification allowed us to largely extend the energetic range evaluated by the parameter, covering wavelengths where the possible differences in the power-law index defining the incident radiation field in AGNs can be properly distinguished. We note that this parameter can also be defined using alternative equivalent transitions for the same ions in a similar spectral range, as in the case for [Ne v] at 24.3 m, and [O iii] at 88 m, which can replace the transitions defined in the above expression.
@@
In the middle panels of Fig. 2, we show the variations of this new parameter as a function of and log . We note that this parameter is only shown when the models predict a certain level of emission for all involved lines, so it cannot be defined for certain low and where the models do not predict any emission of the high-excitation lines. As can be seen, the parameter now clearly traces the hardness of the power-law, as it has lower values for higher values of , consistent with the original optical softness parameter defined by Vílchez & Pagel (1988), which decreases for harder radiation fields. Another important characteristic of this new parameter is its very low dependence on metallicity, given that the assumed AGN SEDs have not got any variation as a function of it, added to the relatively much lower dependence of the mid-IR emission-line emissivities on electron temperature, as compared to optical lines.
@@
On the contrary, one important caveat of this new form of the softness parameter is the fact that the far-IR [O iii], either at 52 m or 88 m, are not covered at low redshift using the available facilities, once the SOFIA observatory has been retired, limiting the number of objects for which an analysis based on this parameter can be carried out. An alternative could be to replace this line using another emission line at a shorter wavelengths, more easily covered by mid-IR observations, corresponding to a similar energy range. One possibility could be the use of the sulphur emission lines of [S iv] at 10.5 m and [S iii] at 18.7 m, covering a very similar energy range (Fig. 1), although with a possible additional uncertainty coming from variations in the S/O abundance ratio (e.g., Pérez-Díaz et al., 2024). In this case, the alternative form of the softness parameter would be
@@
As in the previous case, similar expressions can be explored using the [S iii] line at 33.5 m instead. In the right panels of Fig. 2 we show the dependence of this parameter on and log , where a very similar behavior to that obtained for the version of the parameter based on [O iii] is seen, confirming its potential use to trace the hardness of the ionizing SED in AGNs.
@@
In Fig. 3, we compare the two above-defined parameters represented in two dimensions under the form of diagrams, in order to compare in the same space the observations described in Sect. 2, represented as density points, with the predictions from some of the photoionization models described in Sect. 3. As a first evident impression, the use of the mid-IR lines of [S iii] and [S iv] replacing the far-IR [O iii] results very convenient, given the scarcity of points observed in that regime. In contrast, the IDEOS catalog allowed us to explore a significantly larger sample. As support of this, in our overall sample of 484 objects, we only found 97 presenting a measurement of the far-IR [O iii] line, either at 52 m or 88 m, while [S iii]+[S iv], either with [S iii] at 18.7 m or 33.5 m, can be measured in 480 objects of the compiled sample.
@@
As it can be seen, although the sequences of larger cover the region of the diagram corresponding to high values for both ratios, they consistently show lower values (i.e., the lines of slope 1 in the diagram), while the models with lower are situated in agreement with the expected higher values of . Regarding , the values go in the same direction as already observed in the version of the softness diagram studied for star-forming regions: higher values of are situated in the lower left part of each sequence (i.e., corresponding to low values of Ne23Ne5 and O3O4), while lower values of are situated in the upper right part (i.e., for high values of Ne23Ne5 and O3O4). Both diagrams hence show that they could in principle be used to determine both and .
@@
The equivalent model sequences calculated without dust grains are consistent with studies by several authors who omit them to match the observations of certain UV lines (e.g., Nagao et al. 2006; Dors et al. 2019). Apart from a slight difference at lower values of , this change does not appear to produce differences larger than those caused by other factors, such as geometry (which we confirm in the subsections below). The small differences found in this case are probably motivated by the much lower dependence of the IR lines on dust absorption. This again underlines the advantage of using IR lines for this kind of analysis, as they reduce the uncertainty due to extinction correction in the used emission lines.
4.2 The Hii-Chi-mistry-Teff-IR code for AGNs
The use of the softness parameter based on IR lines, as defined in Eqs. 2 and 3, or using alternative ratios of emission-line fluxes involving low- and high-excitation states of ionization, can serve as a proxy for the hardness of the incident radiation. However, as discussed in Pérez-Montero et al. (2019b) for the optical version of this parameter based on [O ii], [O iii], [S ii], and [S iii], the parameter presents additional dependencies on other functional parameters that can lead to incorrect interpretations for objects presenting different values but also with very different metallicities or excitation levels. A similar behavior was found and studied in Paper I when the corresponding version of the parameter for the mid-IR was examined.
@@ 4.2 The Hii-Chi-mistry-Teff-IR code for AGNs
Similarly, although the mean of the different model sequences correlate with their , there are also deviations as a function of , as can be seen in Fig. 3, and in lesser terms on , as can be seen in Fig. 2, that must be analyzed in order to extract a more precise estimation of the hardness of the radiation field. Thus, the code serves as a tool to minimize these additional dependencies for the analysis of the incident radiation field hardness. Here, we therefore describe the adaptation of the code HCm-Teff-IR so that it includes the emission lines predicted by the photoionization model grids shown in Sect. 3. Additionally, we discuss the appropriate emission-line ratios needed to construct the parameter in its various forms, adapted to the available observational data.
Once the parameters of the models that will be used (i.e., , dust) are chosen, the code first interpolates the tables for the corresponding metallicity, which is treated as an input parameter, expressed as the total oxygen abundance. As in previous versions, if the metallicity is not given as input, the code considers all possible values, implying an enhanced uncertainty. The code then begins an iterative process along all the grid of models to create a -weighted distribution of and log , whose mean and standard deviation are the found solutions. Notice that this is not the classical Bayesian approach in which a unique model with the least resulting is chosen, as this last procedure would be very inefficient to figure out the real uncertainty of the result. The corresponding values are obtained by the code as the sum of the quadratic differences between the observed emission-line ratios and those predicted by the models, which can be calculated with the emission-line fluxes given as input. For this version of the code, we ruled out in the grids those models corresponding to and log , because these do not follow any clear trend useful to derive these parameters from the chosen emission-line ratios, as can be seen in Fig. 2.
@@
The code admits as input the reddening-corrected fluxes in arbitrary units of the emission lines of the ratios involved in the several forms of the parameter, including:444Additionally, other emission lines such as [Ar ii] 7.0 m + [Ar iii] 8.9 m and/or 21.8 m are also accepted by the code when they are given in relation to other high-excitation lines as [Ar v] 7.9 m or 13.1 m. We did not find any object in the compiled samples with these lines, but their interest does not present any doubt considering their potential detection in many targets observed by JWST in the near future.
@@
Specifically for the case of the mid-IR lines in this study, we did not apply any extinction correction due to the relatively low extinction values involved. Additionally, the narrow-line emission in AGNs is produced by an extended and diffuse gas region, which is les affected by the nuclear dust obscuration. The code only considers an emission-line flux for the calculation when it can be used in combination with the emission-line flux for a complementary ionization stage (e.g., [Ne v] is only used if, in addition, [Ne ii] and [Ne iii] are also given). In the case where more than one line is given for a certain ion (e.g., [Ne v], [S iii]), the code uses all forms of the corresponding ratios with equal weight. Therefore, it is up to the user to decide which line(s) to include if all of them are not going to be used.
@@ 5.1 The distribution of for AGNs
First, we calculated the parameters as defined for AGNs in Eqs. 2 and 3 for the compiled sample. The corresponding distributions for these values are represented in the histograms shown in Fig. 4. We note that for these calculations we chose the [O iii] line at 88 m instead of that at 52 m, as this allowed us to calculate the parameter in a larger number of objects in our sample. A similar consideration was taken for the version of the parameter shown in the lower panel, for which we used the [S iii] line at 18.7 m instead of that at 33.5 m. Similarly, we used the [Ne v] line at 14.3 m.
@@
To verify the hypothesis that the studied points can indeed be represented by a double-peaked distribution, we applied a Gaussian mixture model (GMM), assigning each point to a different component. Afterward, to confirm the bimodality found by the GMM, we applied a Shapiro-Wilk normality test independently to each of the two resulting distributions. In each case, the null hypothesis is that the objects do not correspond to a single Gaussian distribution, which is rejected if the resulting p-value is below a 5% threshold. We note that the uncertainties in the line ratios were not taken into account for the normality test. Nevertheless, this would have reduce the weight of uncertain measurements, which typically show larger deviations.
For both representations of the parameter (i.e., using either [O iii] or [S iii]+[S iv]), we verified that all p-values were below the rejection limit, confirming that each one of the two independent distributions found by the GMM are normal and hence that the whole resulting distribution is indeed bimodal. While the overlap region between the two Gaussian components found by the GMM introduces some uncertainty in individual classifications, the statistical trends remain robust as the fraction of galaxies with ambiguous membership is small and does not significantly impact the results.
We also show in Table LABEL:disp_eta the values obtained from this analysis with the mean and standard deviation resulting for each one of the two distributions. The mean values are also shown in Fig. 3. This result could thus indicate that two families of AGNs with a different hardness could co-exist. Nevertheless, given the dependencies shown by both parameters for AGNs (i.e., on and ), a more precise analysis is required to confirm this scenario.
@@ 5.2 Results from HCm-Teff-IR
We applied the HCm-Teff-IR code to our compiled sample of AGNs with measured emission lines in the mid and far-IR spectral range, in order to break the possible degeneracy of the parameter. In particular, we aim to explore whether the observed bimodality is still recovered when we translate the information from the emission lines into the derived parameters and log , the former being a parametrization of the hardness of the incident radiation field in AGNs and the latter representing the excitation state of the gas.
@@
The use of the code also allows for a consistent comparison of results for objects presenting various emission-line combinations across different spectral ranges, thereby increasing the number of objects in the analysis. In this regard, the code admits as input any of the transition lines from [Ne v], [S iii], or [O iii], as described in Sect. 4.2. The only condition to provide a solution is that the Ne23Ne5 emission-line ratio, along with either O3O4 or S34O4, is available.777Additionally, the code admits the emission-line ratio Ar23Ar5, although this is not available for any of the objects compiled for this work.
@@
In Table LABEL:disp we present the results for the objects for which the code finds a solution for , explored in the range from -2.0 to -1.0 and log from -4.0 to -1.0. The table shows the mean values and the standard deviations of the corresponding distributions under different assumptions adopted for the models, including the change in the selected stopping criterion. This value is varied from 2% of free electrons, which in practice implies a totally radiation-bounded geometry, to values 98% and 99.9%, which are those for which most of the sample is covered in the corresponding softness diagrams, as seen in Fig. 3. Results for models computed with or without dust grains mixed with the gas, for each value of , are also given. We did not find any noticeable difference neither in the results nor in their uncertainties for those objects for which a previous determination of the metal content was used by the code, underlining that metallicity is not a crucial factor to derive the hardness or the excitation using IR lines for AGNs.
@@
From an inspection of these average results, it can be seen that the inclusion of dust in the models does not result in a noticeable difference in the outcomes. This is not unexpected, given the very small difference between the model predictions shown by the sequences in Fig. 3. On the contrary, the assumption of different stopping values implies that tends to be higher, on average, for decreasing . This is consistent with the fact that models with = 2% tend to have higher values. However, as discussed in previous sections, this implies that a non-negligible fraction of the studied sample lies outside the model grid, so it is preferable to use the values derived from the code when matter-bounded geometry is assumed, as suggested by previous works (e.g., Meléndez et al., 2011; Pérez-Montero et al., 2019a).
@@
Regarding the differences obtained for the two studied sub-samples (i.e., one for the objects with [O iii] and another for [S iii]+[S iv], which it is essentially the IDEOS sample), there are noticeable differences, particularly for the values, which are significantly higher in the case of IDEOS, with averages lower than -2.0. Nevertheless, Fig. 5 shows that the overall ranges in and log values derived are similar in both cases. Thus, the discrepancy observed could be ascribed to the different sample selection criteria in the two cases. In this regard, objects with [O iii] detections from Herschel/PACS were likely selected among the brightest nearest Seyfert nuclei to guarantee the detection of the faintest far-IR lines. A selection bias toward more luminous objects is likely associated with higher accretion rates.
@@
In addition to the average values, it is also important to analyze the shapes of the parameter distributions to probe how the bimodality found in is interpreted by the code. Figure 5 shows the results obtained for the models with dust grains assuming = 99.9%. The left panel shows the subsample using the O3O4 ratio, either with [O iii] 52m or 88m, while the right panel corresponds to S34O4, using either [S iii] at 18m or 33m. The different symbols in the vs. log diagram correspond to the optical or IR spectral classes indicated in Véron-Cetty & Véron (2010) and Spoon et al. (2022), respectively. See Sect. 5.3 for a discussion on the distribution of different AGN types on this diagram. As in the previous analysis for , the number of objects in this second diagram is much larger (480 against 97), as it involves most of the galaxies in the IDEOS catalog. We note that the total number of objects with a successful determination of and log is 484 (Table LABEL:class), which differs from the number of values in Table LABEL:disp_eta. While is defined only for the 88 m [O iii] line (or alternatively the 18.7 m [S iii] line), the code provides a solution for all the objects with detected [Ne ii], [Ne iii], [Ne v] (14.3 or 24.3 m), [O iv] and either [O iii] (52 or 88 m), or [S iv] and [S iii] (18.7 or 33.5 m). The most important aspect, however, is that the bimodal distribution of the points remains for the two variables derived by the code, that is, for both and log . Besides, the bimodal nature of the two variables appears to be correlated, such that objects with a higher exhibit lower log and vice versa.
@@
Moreover, models help us to realize that the bimodality simultaneously appears for both the radiation field hardness and the excitation. Indeed, the correlation coefficient between these two variables is, on average across the different model assumptions, -0.95. In addition, this correlation can be easily confirmed by quantifying the relative number of objects simultaneously showing a higher – hard radiation (HR) – and a lower log – low excitation (LE) – or vice versa, that is, those with a soft radiation field (SR) and a high excitation (HE), which are the combinations to most of the galaxies apparently belong. As derived from the values shown in Table LABEL:class, this calculation yields that a 91.5% are SRHE or HRLE, confirming that objects with a harder radiation field also have a lower proportion of ionizing photons compared to the number of particles in the gas phase. Taking the weighted mean values obtained in Table LABEL:disp, we can thus define two main AGN families: one with average -1.7 and log -1.5 (SRHE), and another with -1.3 and log -2.4 (HRLE).